Reactive Power Generation has the following capital structure. Its corporate tax rate is 21%.

Security Market Value Required Rate
of Return
Debt $ 20 million 4 %
Preferred stock 10 million 6
Common stock 50 million 10

What is its WACC? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

After-tax return of debt=4*(1-tax rate)

=4*(1-0.21)=3.16%

Total value=(20+10+50)=$80 million

WACC=Respective return*Respective weight

=(20/80*3.16)+(10/80*6)+(50/80*10)

=7.79%